elimination/ addition
3x + y = 7
x + 2y = 4
with steps too

Answer by mananth(16946)

3x+y=7 .............1
x+2y=4 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-6x-2y=-14
x+2y=4
Add the two equations
-5x=-10
/-5
x=2
plug value of x in (1)
3x+y=7
6+y=7
y=7-6
y=1